OpenStudy (anonymous):

As a salesperson at Roaring Waves Beach Supplies, Carissa receives a monthly base pay plus commission on all that she sells. If she sells $500 worth of merchandise in one month, she is paid $405. If she sells $900 of merchandise in one month, she is paid $489. Find Carissa's salary when she sells $1,700 worth of merchandise.

OpenStudy (tkhunny):

Formulas and thinking about it. Pay = Base + Commission Pay = Base + Sales*(Commission Rate) Given information. 405 = Base + 500*(Commission Rate) 489 = Base + 900*(Commission Rate) Solve for missing components.